OpenStudy (anonymous):

Find the diagonal of the rectangular solid with the given measures. l = 18, w = 10, h = 2

OpenStudy (anonymous):

this is tricky give me a moment.

OpenStudy (anonymous):

Okay so you vvant to find the diagonal you're going to need to use something similar to pythagorus' theorem. You knovv\[a^2 + b^2=c^2\]

OpenStudy (anonymous):

except novv its going to be \[a^2+b^2 + c^2=d^2 \]

OpenStudy (anonymous):

a = length b = vvidth c = height